# Wavelet Enhanced Adaptive Frequency Filter for Sequential Recommendation

### v1

@@ B.3 Experimental Settings & Hyperparameters
The following software and hardware environments were used for all experiments: PYTHON 3.9.7, PYTORCH 2.3.0, NUMPY 1.24.3, SCIPY 1.11.1, CUDA 12.2, and NVIDIA Driver 535.104.05, and Intel Xeon Gold 6240 CPU, and 32GB NVIDIA Tesla V100 GPU. For reproducibility, we introduce the best hyperparameter configurations for each dataset in Table 4. We conducted experiments under the following hyperparameters: the $\alpha$ is in {0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9}, and the number of filters $k$ is chosen from {1, 2, 4, 8}. For training, the Adam optimizer is optimized with learning rate in {0.0005, 0.001}.Moreover, to address the sparsity of Amazon datasets and LastFM dataset, a dropout rate of 0.5 is used, compared to 0.1 for MovieLens-1M.
@@ B.4 Model Performance
To investigate the impact of long sequence scenarios on recommendation results, we varied the maximum sequence length $N$ for FMLPRec, BSARec, SLIME4Rec and WEARec. Figure 5 presents the experimental results. We observed that almost all models achieved their best performance at $N=200$, indicating that longer sequence information can more comprehensively represent user behavior patterns. Furthermore, while baseline models showed performance improvements in long-sequence scenarios, they were prone to overfitting, leading to performance convergence. Finally, WEARec consistently outperformed the baselines across all different maximum sequence length settings, and its improvement over baseline models was even more significant in long-sequence scenarios.
@@ C.1 Visualization of Learned $\mathbb{T}$
(a) WEFTRec,WFD,1 layer
(b) WEFTRec,WFD,2 layer

### v2

@@ B.3 Experimental Settings & Hyperparameters
The following software and hardware environments were used for all experiments: PYTHON 3.9.7, PYTORCH 2.3.0, NUMPY 1.24.3, SCIPY 1.11.1, CUDA 12.2, and NVIDIA Driver 535.104.05, and Intel Xeon Gold 6248R CPU, and 40GB NVIDIA Tesla A100 GPU. For reproducibility, we introduce the best hyperparameter configurations for each dataset in Table 4. We conducted experiments under the following hyperparameters: the $\alpha$ is in {0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9}, and the number of filters $k$ is chosen from {1, 2, 4, 8}. For training, the Adam optimizer is optimized with learning rate in {0.0005, 0.001}.Moreover, to address the sparsity of Amazon datasets and LastFM dataset, a dropout rate of 0.5 is used, compared to 0.1 for MovieLens-1M.
@@ B.4 Model Performance
To investigate the impact of long sequence scenarios on recommendation results, we varied the maximum sequence length $N$ for FMLPRec, BSARec, SLIME4Rec and WEARec. Tabel 5 presents the experimental results. We observed that almost all models achieved their best performance at $N=200$, indicating that longer sequence information can more comprehensively represent user behavior patterns. Furthermore, while baseline models showed performance improvements in long-sequence scenarios, they were prone to overfitting, leading to performance convergence. Finally, WEARec consistently outperformed the baselines across all different maximum sequence length settings, and its improvement over baseline models was even more significant in long-sequence scenarios.
@@ C.1 Visualization of Learned $\mathbb{T}$
(a) WEARec,WFE,1 layer
(b) WEARec,WFE,2 layer